OpenStudy (anonymous):

The formula A = 118e^0.024t models the population of a particular city, in thousands, t years after 1998. When will the population of the city reach 140 thousand? A. 2008 B. 2005 C. 2006 D. 2007

OpenStudy (anonymous):

IS B 2005 CORRECT

OpenStudy (anonymous):

140 = 118 e^0.024t 140/118 = e^0.024t ln(140/118) = 0.024t t = ln(140/118)/0.024 = 7 years 1998 + 7 = 2005 so yes

OpenStudy (anonymous):

THANKS